OpenStudy (anonymous):

A health clinic uses a solution of bleach to sterilize petri dishes in which cultures are grown. The sterilization tank contains 150 gal of a solution of 4% ordinary household bleach mixed with pure distilled water. New research indicates that the concentration of bleach should be 8% for complete sterilization. How much of the solution should be drained and replaced with bleach to increase the bleach content to the recommended level?

OpenStudy (ybarrap):

4% of 150 gal has bleach: .04*150 = 6 gallons is bleach We need it to be 8%. That means .08 * 150 gals needs to be bleach: .08*150 = 12 gallons When we drain, what will be left (x) will be 4% bleach, what we add (y) will be 100% bleach. Together we still need it to add up to 150 gallons, so (.04*x + y)/150 = .08 This represents the fraction of fluid that will be bleach after adding y gals of bleach to the fraction of bleach that remains, which is .04x. We also have that x+y=150. So we have 2 equations with 2 unknowns, solvable: $$ \cfrac{.04*x + y}{150} = .08\\ .04*x + y=12\\ x+y=150\\ \implies y=150-x\\ \text{So,}\\ .04x+150-x=12\\ x(.04-1)=12-150\\ x=\cfrac{12-150}{.04-1}\\ x=147.75\\ y=150-x=150-147.75=6.25 $$ So we must drain 6.25 gallons of the original 150 gallons that contained 4% bleach then add back 6.25 gallons of bleach to make a new batch of 150 gallons that now contains 8% bleach. Let's check: $$ \cfrac{6.25+.04*147.75}{150}=0.08 $$ Which is what we wanted. Does this make sense?